OpenStudy (anonymous):

solve each system of equations by the substitution method. Tell whether each system is inconsistent or consistent. If the system is consistent, tell whether it is independent or dependent 2(x+y) -5(x+1)=3 -3x+2y=8

OpenStudy (bahrom7893):

simplify the first equation: \[2x+2y-5x-5=3\]\[-3x+2y=8\]simplify the first equation further: \[2x-5x+2y-5=3\]\[-3x+2y=3+5\]\[-3x+2y=8\]

OpenStudy (bahrom7893):

So now you have:\[-3x+2y=8\]and\[-3x+2y=8\]That's true for all x and y, so this is a consistent, dependent system.

OpenStudy (anonymous):

Thank you! how about this one 5x - 4y=11 3x-y=1

OpenStudy (bahrom7893):

listen i solved like 5 of these for u. Your turn.

OpenStudy (anonymous):

i can watch a lot of videos and still not get it im not good with watching videos and learning that way..

OpenStudy (anonymous):

Just watch the one with patrick... trust me, you'll get something out of it.

OpenStudy (anonymous):

I could say the same thing for when I had to learn Lagrange multipliers, but his stuff helped me.